# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ 2.5 Differential graded Lie groupoids and Lie algebroids
For Ševera differentiation, it is very convenient to work directly in the category $\mathsf{dgMfd}$ of differential graded (dg-)manifolds, because here, the differential is already included in the representation of the functor. [^20] For a discussion of super, i.e. $\mathbbm{Z}_{2}$-graded, Lie groupoids and Lie algebroids, see [Meh06]. In particular, we make the following definition.
@@
Given a split graded manifold $E=(E\rightarrow M)$ with $E$ a graded vector bundle, the grade-shifted tangent bundle $\mathrm{T}[1]E$ is indeed given by the tangent bundle of $E$ with the grading of the tangent fibers shifted appropriately. Moreover, the differential on $\mathrm{T}[1]E$ acts as the sum of the original differential on $E$ and the de Rham differential on $E$, which continues uniquely to all of $\mathrm{T}[1]E$.
To work fully internally to $\mathsf{dgMfd}$, we make the following definition.
@@
A dg-Lie groupoid is a groupoid internal to the category of dg-manifolds $\mathsf{dgMfd}$ with source and target maps surjective submersions. [^21] For the definition of surjective submersions for dg-manifolds, see e.g. [Vys22]. A dg-Lie algebroid is a dg-manifold whose underlying graded ring is non-negatively graded. Together with the evident morphisms, dg-Lie groupoids form the category $\mathsf{dgGrpd}$, and together with their weak and higher morphisms, the 2-category $\mathsf{dg2Grpd}$.
@@
The tangent prolongation dg-Lie algebroid of a dg-Lie algebroid $\mathfrak{g}[1]$ over a dg-manifold $\mathcal{M}$ is the dg-Lie algebroid $\mathrm{T}[1]\mathfrak{g}[1]$ [^22] This graded manifold should be regarded as a Lie 2-algebroid, since a local coordinate description involves coordinate functions in degrees $0,1,2$. It describes the inner derivations of $\mathfrak{g}$. over $\mathrm{T}\mathcal{M}$.
@@
Following [Mac05], we will usually just speak of a tangent groupoid $\mathrm{T}[1]\mathscr{G}$, however, to avoid confusion with the tangent algebroid $\mathrm{T}[1]\mathfrak{g}[1]\to\mathfrak{g}[1]$ we will not omit “prolongation” when speaking of the algebroid structure of $\mathrm{T}[1]\mathfrak{g}[1]$ over $\mathfrak{g}[1]$.
An example of a dg-Lie groupoid important for our discussion is the following.
@@
As one might suspect, Ševera differentiation of Lie groupoids as discussed in Section 2.4 then straightforwardly extends to dg-Lie groupoids.
@@
sending a dg-manifold $\mathscr{M}$ to the set of dg-functors from $\check{\mathscr{C}}(\sigma_{\Theta}(\mathscr{M}))$ to $\mathscr{G}$, where $\mathsf{dgGrpd}$ is the category of dg-Lie groupoids from Definition 2.16.
@@ 2.6 The Weil algebra of a Lie algebroid
An important extension of the Chevalley–Eilenberg algebra $\mathsf{CE}(\mathfrak{g})$ of a Lie algebroid $\mathfrak{g}$ is its Weil algebra, which is the Chevalley–Eilenberg algebra of the tangent prolongation Lie algebroid232323as given in Definition 2.17, for an ordinary Lie algebroid $\mathfrak{g}$ trivially regarded as a dg-Lie algebroid:
@@
The latter is the algebra of functions $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$, together with the differential given by the sum of the Chevalley–Eilenberg and de Rham differentials on $C^{\infty}(\mathfrak{g})$, which uniquely extends to $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$. A more detailed, abstract definition is found in [AAC11], see also [Meh06, JLMP20] for Lie $n$-algebroids. Truncating the latter description to $n=1$, we regard the graded commutative algebra $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$ as the algebra of differential forms $\Omega^{\bullet}(\mathfrak{g}[1])$. As explained above, the Chevalley–Eilenberg differential of $\mathfrak{g}$ can be regarded as a nilquadratic vector field $Q$ of degree $1$ on $\mathfrak{g}[1]$, and the differential in $\mathsf{W}(\mathfrak{g})$ is defined as the sum of the de Rham differential on $\mathfrak{g}[1]$ and the Lie derivative $\mathcal{L}_{Q}$.
@@
$\mathsf{p}_{\rm curv}:\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M\oplus\mathfrak{g}[2]~{}.$ (2.47)
@@
for the underlying graded commutative algebra. The differential $\mathsf{d}_{\mathsf{W}}$ is best described in terms of local coordinates on a chart $U_{M}$ of $M$ over which the bundle $\mathrm{T}[1]\mathfrak{g}[1]$ trivializes. If $m^{a}$ are coordinates on $U_{M}$ and $\bar{m}^{a}$, $\xi^{\alpha}$, and $\tilde{\bar{\xi}}^{\alpha}$ are the coordinates in the fibers [^24] For now we choose a canonical frame, that is, $\bar{m}^{a}$ related to $m^{a}$ in the usual way, and both, $\xi^{a}$ and $\tilde{\bar{\xi}}^{a}$, correspond to the same frame in $\mathfrak{g}$. of $\mathrm{T}[1]M$, $\mathfrak{g}[1]$, and $\mathfrak{g}[2]$, respectively, then the differential reads as
@@
$\mathsf{d}_{\mathsf{W}}\bar{m}^{a}$ $=-\mathtt{r}^{a}_{\alpha}\left(\tilde{\bar{\xi}}^{\alpha}-\omega_{a\beta}^{\alpha}\bar{m}^{a}\xi^{\beta}\right)+\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}\xi^{\alpha}\bar{m}^{b}~{},$
@@
$\left(R^{\rm bas}_{\nabla}\right)^{\alpha}_{\beta\gamma a}$ $=\partial_{a}\mathtt{f}^{\alpha}_{\beta\gamma}+\mathtt{f}^{\delta}_{\beta\gamma}\omega_{a\delta}^{\alpha}+2\omega^{\alpha}_{b[\beta}\partial_{a}\rho^{b}_{\gamma]}-2\omega^{\alpha}_{b[\beta}\omega_{a|\gamma]}^{\delta}\rho^{b}_{\delta}+2\left(\partial_{b}\omega^{\alpha}_{a[\beta}\right)\rho^{b}_{\gamma]}-2\mathtt{f}^{\alpha}_{[\beta|\delta}\omega^{\delta}_{a|\gamma]}$
@@
$\mathsf{d}_{\mathsf{W}}\bar{m}^{a}$ $=-\mathtt{r}^{a}_{\alpha}\left(\bar{\xi}^{\alpha}-\omega_{a\beta}^{\alpha}\bar{m}^{a}\xi^{\beta}-\tfrac{1}{2}\zeta^{\alpha}_{bc}\bar{m}^{b}\bar{m}^{c}\right)+\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}\xi^{\alpha}\bar{m}^{b}~{},$
@@ 3 Local connections on principal groupoid bundles
We follow mostly the ideas of [Car50a, Car50b, KS15a] and in particular [SSS09] and [SS20], and extend them in the evident way to the case of local connections on Lie groupoid bundles.
@@ 3.1 Flat and higher flat connections
where $\nabla$ is any vector bundle connection on $\mathfrak{g}$, $\mathrm{d}^{\nabla}$ denotes the exterior covariant derivative of $\nabla$, and $t_{\nabla_{\rho}}$ is the torsion of $\nabla_{\rho}$, that is,
@@
for all $V,W\in\mathfrak{X}(U_{X})$ and $\mu,\nu\in\Gamma(\mathfrak{g})$, where $\rho$ and [,][-,-][ - , - ] are the anchor and the Lie bracket in $\mathfrak{g}$, respectively; observe, the second equation in (3.2b) is independent of the choice of $\nabla$, in particular, with respect to a frame in $\mathfrak{g}$ and making use of the first equation in (3.2b), the second equation in (3.2b) reads
$\mathrm{d}A^{\alpha}+\frac{1}{2}\mathtt{f}^{\alpha}_{\beta\gamma}A^{\beta}A^{\gamma}=0.$ (3.5)
@@
which encode maps of degree 00, 00, and $-1$, respectively. W.r.t. a frame in $\mathfrak{g}$, the differential $\mathsf{d}$ in $\mathfrak{brst}(U_{X},\mathfrak{g}[1])$ acts on these according to
@@
A coordinate-free version here and in the following is not as straight-forward as before. The major reasons are: many terms have values in $\phi^{*}\mathfrak{g}$ so that a gauge transformation of $\phi$ also affects the values. For example, $A$ projects to $\phi$ (via $\mathfrak{g}\to M$) so that $\delta A$ projects to $\delta\phi$ via $\mathrm{T}\mathfrak{g}\to\mathrm{T}M$; in particular, $\delta A$ is not vertical in $\mathrm{T}\mathfrak{g}$ anymore such that it cannot have values in $\mathfrak{g}$ in general, it has instead values in the tangent prolongation of $\mathfrak{g}$. For technical explanations about this and related terms like the Lie algebra expression of $c$ see [Fis21b, Fis21c].
@@
$c_{M}\in\Omega^{0}(U_{X},\phi^{*}TM)~{},~{}~{}~{}c_{\mathfrak{g}}\in\Omega^{0}(U_{X},\phi^{*}\mathfrak{g})~{},~{}~{}~{}\lambda\in\Omega^{1}(U_{X},\phi^{*}\mathfrak{g})~{},$ (3.12)
@@ 3.2 Non-flat connections as higher flat connections
For higher gauge theory, the appropriate reduction was identified in [SSS09, Def. 34] as partially flat homotopies, i.e. homotopies for which the auxiliary potentials taking the role of the curvatures vanish in the homotopy directions. This ensures that gauge transformations map flat potentials to flat potentials.
@@
$\mathrm{d}\lambda^{\alpha}$ $=-\mathrm{d}\chi^{\alpha}-\left(\mathtt{f}^{\alpha}_{\beta\gamma}+\mathtt{r}^{a}_{\gamma}\omega^{\alpha}_{a\beta}\right)\left(A_{\mathfrak{g}}^{\beta}\chi^{\gamma}+c^{\beta}_{\mathfrak{g}}\lambda^{\gamma}\right)-\left(\omega^{\alpha}_{a\beta}-\zeta^{\alpha}_{ab}\mathtt{r}^{b}_{\beta}\right)\left(A_{M}^{a}\chi^{\beta}+c^{a}_{M}\lambda^{\beta}\right)$ (3.17)
$~{}~{}~{}~{}-\left(R_{\nabla}^{\rm bas}\right)^{\alpha}_{\beta\gamma a}\left(c^{\beta}_{\mathfrak{g}}A^{\gamma}_{\mathfrak{g}}c^{a}_{M}-\frac{1}{2}c^{\beta}_{\mathfrak{g}}c^{\gamma}_{\mathfrak{g}}A^{a}_{M}\right)+3\left(\frac{1}{6}\left(\mathrm{d}^{\nabla}\right)^{\alpha}_{abc}-\frac{1}{2}\zeta^{\alpha}_{ad}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{bc}\right)c^{a}_{M}c^{b}_{M}A_{M}^{c}$
$~{}~{}~{}~{}-\left(R_{\nabla}+\nabla^{\rm bas}\zeta\right)^{\alpha}_{ab\beta}\left(c_{M}^{a}A_{M}^{b}c_{\mathfrak{g}}^{\beta}-\frac{1}{2}c^{a}_{M}c^{b}_{M}A^{\beta}_{\mathfrak{g}}\right)~{},$
@@
There are now two solutions to this. Either, we impose the condition that $A_{M}^{a}=0$ for all the connections we consider. This is the Lie algebroid analogue of the fake curvature condition of higher gauge theory, cf. [SS20] and [BJFJ+24] for a discussion. In the case of higher gauge theory, this has already been shown to be too restrictive for many physical applications. Alternatively, we can demand that $\nabla$ is chosen such that $R_{\nabla}^{\rm bas}=0$. This is the Lie algebroid analogue of the notion of adjustment as defined in [SS20]. This perspective proved to be the right one for important physical applications [BJFJ+24], and it also leads to the correct differential refinement of string structures [SS20, KS20, RSW22] and [TD23]. In the context of curved gauge theories it was also discovered that the closure of the commutator of infinitesimal gauge transformations requires such a condition, see [BKS05, MS09, Fis21b, Fis21c]. We therefore follow this approach.
@@
$R_{\nabla}^{\rm bas}=0~{},~{}~{}~{}R_{\nabla}+\nabla^{\rm bas}\zeta=0~{},~{}~{}~{}\mbox{and}~{}~{}~{}\tfrac{1}{6}\left(\mathrm{d}^{\nabla}\zeta\right)^{\alpha}_{abc}-\tfrac{1}{2}\zeta^{\alpha}_{ad}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{bc}=0~{},$ (3.20)
where $\mathrm{d}^{\nabla}$ is the common notation for the exterior covariant derivative w.r.t. $\nabla$; as a reminder, in this case we have
@@
As pointed out in [Fis21a, Fis21b], this means that the first two equations are of a cohomological type; integrated in [Fis22] (in the case of Lie group bundles as structure); however, this requires introducing exterior covariant derivatives of Lie algebroid connections and a double-degree version of forms.

### v3

@@ 2.5 Differential graded Lie groupoids and Lie algebroids
For Ševera differentiation, it is very convenient to work directly in the category $\mathsf{dgMfd}$ of differential graded (dg-)manifolds, because here, the differential is already included in the representation of the functor. [^21] For a discussion of super, i.e. $\mathbbm{Z}_{2}$-graded, Lie groupoids and Lie algebroids, see [Meh06]. We briefly develop this picture in the following.
@@
Given a split dg-manifold $E=(E\rightarrow M)$, i.e. a graded vector bundle $E$ over a manifold $M$, the grade-shifted tangent bundle $\mathrm{T}[1]E$ is indeed given by the tangent bundle of $E$ with the grading of the tangent fibers shifted appropriately. Moreover, the differential on $\mathrm{T}[1]E$ acts as the sum of the original differential on $E$ and the de Rham differential on $E$, which continues uniquely to all of $\mathrm{T}[1]E$.
To work fully internally to the category $\mathsf{dgMfd}$ of differential graded manifolds, we need the following definition.
@@
A dg-Lie groupoid is a groupoid internal to the category of dg-manifolds $\mathsf{dgMfd}$ with source and target maps surjective submersions. [^22] For the definition of surjective submersions for dg-manifolds, see e.g. [Vys22]. A dg-Lie algebroid is a dg-vector bundle (in the evident sense) with fibers graded vector spaces concentrated in non-negative degrees. Together with the evident morphisms, dg-Lie groupoids form the category $\mathsf{dgGrpd}$, and together with their weak and higher morphisms, the 2-category $\mathsf{dg2Grpd}$.
@@
The tangent prolongation dg-Lie algebroid of a dg-Lie algebroid $\mathfrak{g}[1]$ over a dg-manifold $\mathcal{M}$ is the dg-Lie algebroid [^23] This graded manifold should be regarded as a Lie 2-algebroid, since a local coordinate description involves coordinate functions in degrees $0,1,2$. It describes the inner derivations of $\mathfrak{g}$. $\mathrm{T}[1]\mathfrak{g}[1]$ over $\mathrm{T}[1]\mathcal{M}$.
@@
Following [Mac05], we will usually just speak of a tangent groupoid $\mathrm{T}[1]\mathscr{G}$, however, to avoid confusion with the tangent algebroid $\mathrm{T}[1]\mathfrak{g}[1]\to\mathfrak{g}[1]$ we will not omit “prolongation” when speaking of the algebroid structure of $\mathrm{T}[1]\mathfrak{g}[1]$ over $\mathrm{T}[1]\mathcal{M}$.
An example of a dg-Lie groupoid which is important for our discussion is the following.
@@
As one would expect, Ševera differentiation of Lie groupoids as discussed in Section 2.4 then straightforwardly extends to $\mathsf{dgGrpd}$, the category of dg-Lie groupoids from Definition 2.17.
@@
sending a dg-manifold $\mathscr{M}$ to the set of dg-functors from $\check{\mathscr{C}}(\sigma_{\Theta}(\mathscr{M}))$ to $\mathscr{G}$.
@@
Given such a right action, we can construct the corresponding action groupoid, cf., e.g., [Mac05, Def. 1.6.10] for the case of ordinary groupoids.
Definition 2.25 (Action groupoid).
Consider a right action $(\mathbin{\vartriangleleft},\psi)$ of a dg-Lie groupoid $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ on a dg-manifold $N$. The corresponding action groupoid is given by
with structure maps
for $g_{1,2}\in\mathscr{G}$ and $n\in N$ with $\psi(n)=\mathsf{t}(g)$ and $\mathsf{s}(g_{1})=\mathsf{t}(g_{2})$.
Similarly, one can construct the action groupoid for left actions $(\vartriangleright,\psi)$. Infinitesimally, an action groupoid differentiates to an action Lie algebroid.
@@ 2.6 The Weil algebra of a Lie algebroid
An important extension of the Chevalley–Eilenberg algebra $\mathsf{CE}(\mathfrak{g})$ of a Lie algebroid $\mathfrak{g}$ is its Weil algebra $\mathsf{W}(\mathfrak{g})$, which is the Chevalley–Eilenberg algebra of the tangent prolongation Lie algebroid242424as given in Definition 2.18, for an ordinary Lie algebroid $\mathfrak{g}$ trivially regarded as a dg-Lie algebroid:
@@
The latter is the algebra of functions $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$, together with the differential $\mathsf{d}_{\mathsf{W}}$ given by the sum of the Chevalley–Eilenberg and de Rham differentials on $C^{\infty}(\mathfrak{g})$, which uniquely extends to $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$. Explicitly, we have
$\mathsf{d}_{\mathsf{W}}f$ $=\mathrm{d}f\circ\rho+\mathrm{d}f~{},$ (2.48)
$\mathsf{d}_{\mathsf{W}}\omega$ $=\mathsf{d}_{\mathsf{CE}}\omega+\nabla^{*}\omega+\omega[1]~{},$
$\mathsf{d}_{\mathsf{W}}(\mathrm{d}f)$ $=-\left(\mathrm{d}f\circ\rho\right)[1]-\nabla^{*}\left(\mathrm{d}f\circ\rho\right)~{},$
$\mathsf{d}_{\mathsf{W}}\omega[1]$ $=-\mathsf{d}_{\mathsf{W}}(\mathrm{d}_{\mathsf{CE}}\omega+\nabla^{*}\omega)~{}$
for all $f\in C^{\infty}(M)$, $\omega\in\Gamma(\mathfrak{g}[1]^{*})$, where $[1]$ is the degree shift $\Gamma(\mathfrak{g}[1]^{*})\to\Gamma(\mathfrak{g}[2]^{*})$ and $\nabla^{*}$ is the induced connection on the dual vector bundle defined by
$\left(\nabla_{V}^{*}\omega\right)(\alpha)=-\omega(\nabla_{V}\alpha)+V(\omega(\alpha))~{},$ (2.49)
for vector fields $V$ and $\alpha\in\Gamma(\mathfrak{g})$.
A more detailed discussion is found in [AAC11], see also [Meh06, JLMP20] for Lie $n$-algebroids. Truncating the latter description to $n=1$, we regard the graded commutative algebra $C^{\infty}(\mathrm{T}[1]\mathfrak{g}[1])$ as the algebra of differential forms $\Omega^{\bullet}(\mathfrak{g}[1])$. As explained above, the Chevalley–Eilenberg differential of $\mathfrak{g}$ can be regarded as a nilquadratic vector field $Q$ of degree $1$ on $\mathfrak{g}[1]$, and the differential in $\mathsf{W}(\mathfrak{g})$ is defined as the sum of the de Rham differential on $\mathfrak{g}[1]$ and the Lie derivative $\mathcal{L}_{Q}$.
@@
$\mathsf{p}_{\rm curv}\colon\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M\oplus\mathfrak{g}[2]~{}.$ (2.51)
@@
for the underlying graded commutative algebra.
It is convenient to describe the differential $\mathsf{d}_{\mathsf{W}}$ in terms of local coordinates on a chart $U_{M}$ of $M$ over which the bundle $\mathrm{T}[1]\mathfrak{g}[1]$ trivializes. If $m^{a}$ are coordinates on $U_{M}$ and $\bar{m}^{a}$, $\xi^{\alpha}$, and $\tilde{\bar{\xi}}^{\alpha}$ are the coordinates in the fibers [^25] For now we choose a canonical frame, that is, $\bar{m}^{a}$ related to $m^{a}$ in the usual way, and both, $\xi^{a}$ and $\tilde{\bar{\xi}}^{a}$, correspond to the same frame in $\mathfrak{g}$. of $\mathrm{T}[1]M$, $\mathfrak{g}[1]$, and $\mathfrak{g}[2]$, respectively, then the differential reads as
@@
$\mathsf{d}_{\mathsf{W}}\bar{m}^{a}$ $=-\mathtt{r}^{a}_{\alpha}\left(\tilde{\bar{\xi}}^{\alpha}-\omega_{b\beta}^{\alpha}\bar{m}^{b}\xi^{\beta}\right)+\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}\xi^{\alpha}\bar{m}^{b}~{},$
@@
$\left(R^{\rm bas}_{\nabla}\right)^{\alpha}_{\beta\gamma a}$ $=\partial_{a}\mathtt{f}^{\alpha}_{\beta\gamma}+\mathtt{f}^{\delta}_{\beta\gamma}\omega_{a\delta}^{\alpha}+2\omega^{\alpha}_{b[\beta}\partial_{a}\mathtt{r}^{b}_{\gamma]}-2\omega^{\alpha}_{b[\beta}\omega_{a|\gamma]}^{\delta}\mathtt{r}^{b}_{\delta}+2\left(\partial_{b}\omega^{\alpha}_{a[\beta}\right)\mathtt{r}^{b}_{\gamma]}-2\mathtt{f}^{\alpha}_{[\beta|\delta}\omega^{\delta}_{a|\gamma]}$
@@
$\mathsf{d}_{\mathsf{W}}\bar{m}^{a}$ $=-\mathtt{r}^{a}_{\alpha}\left(\bar{\xi}^{\alpha}-\omega_{b\beta}^{\alpha}\bar{m}^{b}\xi^{\beta}-\tfrac{1}{2}\zeta^{\alpha}_{bc}\bar{m}^{b}\bar{m}^{c}\right)+\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}\xi^{\alpha}\bar{m}^{b}~{},$
@@ 3 Local connections on principal groupoid bundles
We follow mostly the ideas of [Car50a, Car50b, KS15a] and in particular [SSS09] and [SS20], and extend them in the evident way to the case of local connections on Lie groupoid bundles. As a new result, we show that adjustments amount to Cartan connections.
@@ 3.1 Flat and higher flat connections
where $\nabla$ is any vector bundle connection on $\mathfrak{g}$, $\mathrm{d}^{\nabla}$ denotes the exterior covariant derivative of $\nabla$, and $t_{\nabla_{\rho}}$ is the torsion of $\nabla_{\rho}$. That is,
@@
for all $V,W\in\mathfrak{X}(U_{X})$ and $\mu,\nu\in\Gamma(\mathfrak{g})$, where $\rho$ and [,][-,-][ - , - ] are the anchor and the Lie bracket in $\mathfrak{g}$, respectively; observe, the second equation in (3.2b) is independent of the choice of $\nabla$, in particular, with respect to a frame given by a basis $(e_{\alpha})$ in $\mathfrak{g}$ and making use of the first equation in (3.2b), the second equation in (3.2b) reads
$\mathrm{d}A^{\alpha}+\frac{1}{2}\mathtt{f}^{\alpha}_{\beta\gamma}A^{\beta}A^{\gamma}=0~{}.$ (3.5)
@@
which encode maps of degree 00, 00, and $-1$, respectively. With respect to a frame $(e_{\alpha})$ in $\mathfrak{g}$, the differential $\mathsf{d}$ in $\mathfrak{brst}(U_{X},\mathfrak{g}[1])$ acts on these according to
@@
Here and in the following, a coordinate-free version is not as straight-forward as before: many terms have values in $\phi^{*}\mathfrak{g}$ so that a gauge transformation of $\phi$ have a much wider effect than in ordinary gauge theory. For example, the one-form potential $A$ projects to $\phi$ (via $\mathfrak{g}\to M$) so that $\delta A$ projects to $\delta\phi$ via $\mathrm{T}\mathfrak{g}\to\mathrm{T}M$; in particular, $\delta A$ is not vertical in $\mathrm{T}\mathfrak{g}$ anymore such that it cannot have values in $\mathfrak{g}$ in general, it has instead values in the tangent prolongation of $\mathfrak{g}$. For technical explanations of this and related terms like the Lie algebra expression of $c$ see [Fis21b, Fis21c].
@@
$c_{M}\in\Omega^{0}(U_{X},\phi^{*}\mathrm{T}M)~{},~{}~{}~{}c_{\mathfrak{g}}\in\Omega^{0}(U_{X},\phi^{*}\mathfrak{g})~{},~{}~{}~{}\lambda\in\Omega^{1}(U_{X},\phi^{*}\mathfrak{g})~{},$ (3.12)
@@ 3.2 Non-flat connections as higher flat connections
For higher gauge theory, the appropriate reduction was identified in [SSS09, Def. 34] as partially flat homotopies, i.e. homotopies for which the auxiliary potentials taking the role of the curvatures vanish in the homotopy directions. This ensures that gauge transformations map flat potentials to flat potentials. In other words, the curvature analogue of gauge transformations are zero or gauge transformations are flat.
@@
$\mathsf{d}\underline{\lambda}^{\alpha}$ $=-\mathrm{d}\underline{\chi}^{\alpha}-\left(\mathtt{f}^{\alpha}_{\beta\gamma}+\mathtt{r}^{a}_{\gamma}\underline{\omega}^{\alpha}_{a\beta}\right)\left(\underline{A}_{\mathfrak{g}}^{\beta}\underline{\chi}^{\gamma}+\underline{c}^{\beta}_{\mathfrak{g}}\underline{\lambda}^{\gamma}\right)-\left(\underline{\omega}^{\alpha}_{a\beta}-\zeta^{\alpha}_{ab}\mathtt{r}^{b}_{\beta}\right)\left(\underline{A}_{M}^{a}\underline{\chi}^{\beta}+\underline{c}^{a}_{M}\underline{\lambda}^{\beta}\right)$ (3.17)
$~{}~{}~{}~{}-\left(R_{\nabla}^{\rm bas}\right)^{\alpha}_{\beta\gamma a}\left(\underline{c}^{\beta}_{\mathfrak{g}}\underline{A}^{\gamma}_{\mathfrak{g}}\underline{c}^{a}_{M}-\frac{1}{2}\underline{c}^{\beta}_{\mathfrak{g}}\underline{c}^{\gamma}_{\mathfrak{g}}\underline{A}^{a}_{M}\right)+3\left(\frac{1}{6}\left(\mathrm{d}^{\nabla}\zeta\right)^{\alpha}_{abc}-\frac{1}{2}\zeta^{\alpha}_{ad}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{bc}\right)\underline{c}^{a}_{M}\underline{c}^{b}_{M}\underline{A}_{M}^{c}$
$~{}~{}~{}~{}-\left(R_{\nabla}+\nabla^{\rm bas}\zeta\right)^{\alpha}_{ab\beta}\left(\underline{c}_{M}^{a}\underline{A}_{M}^{b}c_{\mathfrak{g}}^{\beta}-\frac{1}{2}\underline{c}^{a}_{M}\underline{c}^{b}_{M}\underline{A}^{\beta}_{\mathfrak{g}}\right)~{},$
@@
There are now two solutions to this. Either, we impose the condition that $A_{M}^{a}=0$ for all the connections we consider. This is the Lie algebroid analogue of the fake curvature condition of higher gauge theory, cf. [SS20] and [BJFJ+25] for a discussion. In the case of higher gauge theory, this has already been shown to be too restrictive for many physical applications. Alternatively, we can demand that $\nabla$ is chosen such that $R_{\nabla}^{\rm bas}=0$. This is the Lie algebroid analogue of the notion of adjustment as defined in [SS20]. This perspective proved to be the right one for important physical applications [BJFJ+25], and it also leads to the correct differential refinement of string structures [SS20, KS20, RSW22] and [TD23]. In the context of curved gauge theories it was also discovered that the closure of the commutator of infinitesimal gauge transformations requires such a condition, see [BKS05, MS09, Fis21b, Fis21c]. We therefore follow this approach.
@@
$R_{\nabla}^{\rm bas}=0~{},~{}~{}~{}R_{\nabla}+\nabla^{\rm bas}\zeta=0~{},~{}~{}~{}\mbox{and}~{}~{}~{}\tfrac{1}{6}\left(\mathrm{d}^{\nabla}\zeta\right)^{\alpha}_{abc}-\tfrac{1}{6}\zeta^{\alpha}_{[a|d}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{|bc]}=0~{},$ (3.20)
where we anti-symmetrize over $a$, $b$, and $c$ in the second summand of the third equation, and $\mathrm{d}^{\nabla}$ is the common notation for the exterior covariant derivative w.r.t. $\nabla$; as a reminder, in this case we have
@@
As pointed out in [Fis21a, Fis21b], this means that the first two equations are of a cohomological type; integrated in [Fis22] (in the case of Lie group bundles as structure). However, this requires introducing exterior covariant derivatives of Lie algebroid connections and a double-degree version of forms.